OpenStudy (anonymous):

Find the y-intercept of the exponential function f(x) = -2(3)^x.

OpenStudy (danjs):

A quick way to find the Y intercept, is let X=0,

OpenStudy (danjs):

\[y = f(0) = -2* 3^0 = -2*1 = -2\] When x is zero, y is -2 The y intercept is -2